OpenStudy (smart8thgrader):

Can someone help me with 5 questions please????? I WILL MEDAL AND FAN!!!!!

OpenStudy (smart8thgrader):

1Determine which two values the following expression is between. https://media.studyisland.com/cgi-bin/mimetex.cgi?9 \sqrt[3]{29} 27.7 and 27.8 27.6 and 27.7 27.5 and 27.6 27.4 and 27.5.

OpenStudy (smart8thgrader):

@pooja195

pooja195 (pooja195):

is this a test?

OpenStudy (smart8thgrader):

if you count studyisland than yes

OpenStudy (misty1212):

study island lol

OpenStudy (misty1212):

HI!

OpenStudy (smart8thgrader):

yea and iam really bad at it LOL

OpenStudy (misty1212):

link does not work!

OpenStudy (misty1212):

i am reading on your question \[\sqrt[3]{29}\]which you can do nothing with is there another part?

OpenStudy (misty1212):

is it \[9\times \sqrt[3]{29}\]?

OpenStudy (smart8thgrader):

no iwill give you the equation

OpenStudy (smart8thgrader):

https://media.studyisland.com/cgi-bin/mimetex.cgi?9 \sqrt[3]{29}

OpenStudy (smart8thgrader):

@misty1212

OpenStudy (misty1212):

really it does not work try it and you will see

OpenStudy (smart8thgrader):

i know can i just give you the next question?

OpenStudy (misty1212):

i got it

OpenStudy (misty1212):

here is your answer http://www.wolframalpha.com/input/?i=9*%5Csqrt%5B3%5D%7B29%7D

OpenStudy (misty1212):

it is between \(27.6\) and \(27.7\)

OpenStudy (smart8thgrader):

thanks now for number 2

OpenStudy (smart8thgrader):

Determine which two values the following expression is between. https://media.studyisland.com/cgi-bin/mimetex.cgi?4 \sqrt[3]{7} 8 and 9 9 and 10 10 and 11 7 and 8

OpenStudy (smart8thgrader):

@misty1212

OpenStudy (smart8thgrader):

BYE BYE

OpenStudy (smart8thgrader):

@misty1212

OpenStudy (misty1212):

\[4\sqrt[3]{7}\]?

OpenStudy (misty1212):

type it in to wolfram and see what you get like this http://www.wolframalpha.com/input/?i=4*%5Csqrt%5B3%5D%7B7%7D